# Undesirable Memorization in Large Language Models: A Survey

### v2

@@ Abstract
While recent research increasingly showcases the remarkable capabilities of Large Language Models (LLMs), it is equally crucial to examine their associated risks. Among these, privacy and security vulnerabilities are particularly concerning, posing significant ethical and legal challenges. At the heart of these vulnerabilities stands memorization, which refers to a model’s tendency to store and reproduce phrases from its training data. This phenomenon has been shown to be a fundamental source to various privacy and security attacks against LLMs.
In this paper, we provide a taxonomy of the literature on LLM memorization, exploring it across three dimensions: granularity, retrievability, and desirability. Next, we discuss the metrics and methods used to quantify memorization, followed by an analysis of the causes and factors that contribute to memorization phenomenon. We then explore strategies that are used so far to mitigate the undesirable aspects of this phenomenon. We conclude our survey by identifying potential research topics for the near future, including methods to balance privacy and performance, and the analysis of memorization in specific LLM contexts such as conversational agents, retrieval-augmented generation, and diffusion language models.
Given the rapid research pace in this field, we also maintain a dedicated repository of the references discussed in this survey [^1] https://github.com/alistvt/undesirable-llm-memorization, which will be regularly updated to reflect the latest developments.
@@ I Introduction
Despite these impressive advancements of LLMs, researchers have shown that there are certain problems with these models, including hallucination [9], bias [10], and privacy and security vulnerabilities [11]. In the context of data privacy, memorization is one of the core sources of concern. Memorization in LLM refers to the model’s tendency to store and reproduce exact phrases or passages from the training data rather than generating novel or generalized outputs. While memorization can be advantageous in knowledge-intensive benchmarks, such as factual recall tasks or domain-specific question answering [12], it also introduces ethical and legal challenges: models may inadvertently reveal sensitive or private information included in their training data, posing significant privacy and security risks [13, 14]. In addition, the ability of LLMs to repeat verbatim copyrighted or proprietary text from their training data raises issues related to intellectual property infringement [15, 16].
These challenges motivate the need to further explore memorization in LLMs to effectively tackle the associated challenges. In this paper, we provide an overview of aspects related to memorization, emphasizing the need for further exploration of this topic, thereby balancing model performance with the risks of privacy breaches and ethical concerns.
@@ I-B Data selection
The selection of works included in this survey was guided by the goal of capturing a comprehensive view of the literature surrounding memorization in LLMs, with a particular focus on its undesirable aspects. However, to provide a well-rounded taxonomy and to contextualize these challenges, we also consider works that explore the desirable aspects of memorization, such as generalization and knowledge retention, as they offer complementary insights into the mechanisms at play.
@@
The process of identifying relevant papers was iterative, beginning with the most widely cited early studies by recognized researchers in the field (14 papers). These foundational and influential works were selected as they represent key milestones that have shaped current understandings of memorization in LLMs. Then we included the papers drawn from keyword searches in article repositories. We used Google Scholar [^2] https://https://scholar.google.com/ and arXiv [^3] https://arxiv.org/, as they are predominantly used by researchers in the computer science and computational linguistics domains [22]. For Google Scholar we restricted our keyword search to the papers that include the term “Memorization” in their title and include “Language Model” in their body (memorization is also used as a term in biology). For the scanning of arXiv, we collected the papers that included both of the words “Memorization” and “Privacy” in their abstract. For inclusion, we selected articles published before January 2025.
@@
Figure 1: Our data selection process: Starting with 14 foundational papers as the seed, our literature search expanded through keyword search in Google Scholar (223 papers) and arXiv (111 papers), followed by manual filtering, reference chaining, and exploration of adjacent areas, concluding in a final selection of 99 papers.
@@ I-C Paper organization
Figure 2: Graphic summary of our survey
@@ II Spectrum of Memorization
In this section, we provide an overview of the available definitions for memorization in various dimensions and categorize the existing work according to these dimensions. The overall summary of our findings is presented in Table III along with the other aspects.
@@ II-A Granularity of memorization
This definition is more strict and does not capture the memorization under different forms of sampling methods. It rather considers only the greedy sampling [see Section IV-E].
@@
Throughout the rest of this paper, we use the word memorization interchangeably with the verbatim memorization, as it is the one that is mostly associated with the undesirable aspects of memorization.444discussed in section II-C
@@ II-B Retrievability
(Extractable Memorization) Let $LM(p)$ denote the output of a language model when prompted with $p$. An example $s$ from the training set $S$ is extractably memorized if an adversary (without access to $S$) can construct a prompt $p$ that makes the model produce $s$ (that is, $LM(p)=s$).
@@
(Discoverable memorization) Let $LM(p)$ denote the output of a language model when prompted with $p$. For a context $c=(p,s)$ from the training set $S$, we say that $s$ is discoverably memorized if $LM(p)=s$.
@@
(k-discoverable memorization) For an example $c=(p,s)$ from the training set, string $s$ is said to be k-discoverable if $s$ is discoverable and $p$ is consisted of $k$ tokens [42].
@@ II-C Desirability
  • Copyright violation: Memorization of copyrighted material may lead to legal challenges, especially if the model can reproduce substantial portions of protected works [16, 50, 51].
@@
  • Transparency and auditing: Memorization poses challenges for transparency and auditing by making it difficult to decide whether specific outputs stem from generalization or memorized content, leading to accountability and interpretability issues in LLM systems.
@@
  • Knowledge retention: LLMs are often deliberately designed to memorize facts, concepts, and general knowledge to enhance their performance in tasks like question answering and information retrieval [55, 56].
@@
  • Alignment goals: Deliberate memorization can be utilized in the AI alignment phase to inject desired behaviors and values into the model [57].
@@ III-A String match
Measuring verbatim memorization (Section Definition) with the methods discussed in the retrievability section (Section II-B), requires exhaustively interacting with the model by inputting different prompts and comparing the model output to the training data. Then the attack success rate is measured by dividing the portion of memorized text by the size of the training data. Since there are infinite combinations of tokens that one can feed the model, this method usually falls short in providing the accurate amount of memorization, rather it can provide a good approximation on memorization lower bound. Since this is directly linked to the memorization risks, it is predominantly used by the researchers [47, 40]. As the exact match metric is too sensitive to small perturbations in the generation, approximate match (Section Definition) has been introduced [26] to capture more memorized samples by bypassing negligible changes in the outputs.
@@ III-B Exposure metric
Definition (Exposure).
Given a canary $s[r]$, a model with parameters $\theta$, and the randomness space R, the exposure of $s[r]$ is [47]:
@@ III-C Inference attacks
These attacks aim to determine whether a specific data point was part of the model’s training set. Shokri et al. [59] introduced this concept for machine learning models, and it has since been adapted for language models. Doing MIA on LLMs typically involves determining the model’s confidence in a given text and using it as an indicator of whether the text was part of the training data. This is often done by computing the perplexity of the sequence [60, 61, 62, 63, 64, 65, 66].
@@
These attacks attempt to extract specific pieces of information from the model that were present in its training data. Different works have demonstrated the feasibility of such attacks on language models, showing that private information could be extracted through prompting with different strategies [67, 40, 68, 47, 69, 66].
@@ III-D Counterfactuality
Previous work analyzed the memorization of large language models on sensitive information (e.g. phone numbers) in the training data [13] or synthetically injected ‘canaries’ [24, 70]. However, not all the memorized texts are equally interesting. Zhang et al. [71] propose another measure of memorization which is counterfactual memorization. The idea is to see how the presence or absence of a sample of the dataset, affects the performance of the model on the same sample. This measure has similarities with the definition of differential privacy. [^7] Differential privacy as a mitigation strategy will be discussed in Section V-B.
@@
III-E Prompt Compression
Schwarzschild et al. [72] introduce the Adversarial Compression Ratio (ACR) as a novel metric to assess memorization in large language models (LLMs). ACR evaluates whether a string from the training data is memorized by determining if it can be elicited with a significantly shorter adversarial prompt. The method uses the GCG algorithm [73] to find the most compressed prompt. This approach aligns conceptually with Kolmogorov complexity [^8] Kolmogorov complexity measures the shortest possible set of instructions (or program) needed for a computer (Turing machine) to generate a specific output., as it measures the minimum description length of a string but adapts the concept for practical use in LLMs by focusing on adversarial prompts. ACR provides a flexible and compute-efficient framework to evaluate memorization for arbitrary strings, particularly in scenarios like monitoring unlearning and ensuring compliance with data usage policies. This metric offers practical and legal insights into potential misuse of training data, addressing critical concerns in LLM development.
@@ IV Influencing factors and dynamics of memorization
Understanding the factors that influence memorization in LLMs is crucial for developing more efficient, secure, and privacy-preserving systems. This section explores various aspects that affect memorization, from model architecture to training processes.
IV-A Model capacity
@@
IV-B Training data characteristics
The nature of the training data heavily influences memorization. Lee et al. [77] develop tools to deduplicate training data and show that models trained on deduplicated data would produce memorized text ten times less frequently. Kandpal et al. [23] show that a sequence appearing 10 times in the training data is, on average, generated approximately 1000 times more frequently than a sequence that appears only once. A study done by Tirumala et al. [25] on memorization of different parts of speech, reveals that nouns and numbers are memorized significantly faster than other parts of speech, likely because they serve as unique identifiers for specific samples. Prashanth et al. [78] categorize memorization into recitation (memorization of highly duplicated sequences), reconstruction (generation using learned templates or patterns), and recollection (memorization of rare, non-template sequences). They argue that memorization is a multi-faceted phenomenon influenced by duplication, template patterns, and rarity, requiring nuanced analysis.
As simple sequences, such as repeated patterns or numbers, are easily memorized by models but often lack substantive content or sensitive information, distinguishing the memorization of these trivial sequences from more complex and meaningful ones would be essential. Duan et al. [28] observed that data with lower z-complexity [^9] Z-complexity is a way to measure how much temporary memory (workspace) a computer needs to generate a specific output using the shortest possible instructions. Unlike Kolmogorov complexity, which only looks at the length of the instructions, Z-complexity also considers the memory used during the process. leads to faster decreases in training loss, as more compressible patterns are memorized more quickly. They further demonstrated that strings of varying complexity exhibit distinct memorization curves, with lower-complexity strings being memorized more easily even for smaller repeats, following a log-linear relationship in memorization probability.
IV-C Input and prompting strategies
Carlini et al. [13], McCoy et al. [79], Kandpal et al. [23] show that longer prompts increase the likelihood of triggering memorized sequences, making it easier for language models to regurgitate training data. Moreover, methods like prefix tuning [80] and prompt engineering have been employed to maximize memorization. Ozdayi et al. [81] introduce a novel approach using prompt tuning to control memorized content extraction rates in LLMs. Wang et al. [82] introduce a dynamic, prefix-dependent soft prompt approach to elicit more memorization. By generating soft prompts based on input variations, this method outperforms previous techniques in extracting memorized content across both text and code generation tasks. Kassem et al. [83] investigate how instruction-tuning influences memorization in LLMs, showing that instruction-tuned models can reveal as much or more training data as base models. They propose a black-box prompt optimization method where an attacker LLM generates instruction-based prompts, achieving higher memorization levels than direct prompting approaches. Weller et al. [84] propose ‘according-to’ prompting, a technique that directs LLMs to ground responses in previously observed text.
IV-D Tokenization
@@
Kharitonov et al. [85] explore the impact of the tokenizer on memorization. They experiment with the size of the sub-word vocabulary learned through Byte-Pair Encoding (BPE) and demonstrate that increasing the sub-word vocabulary significantly affects the model’s ability and inclination to memorize training data. Furthermore, models with larger vocabularies are more likely to reproduce training data when given specific prompts. The authors suggest that this effect stems from the reduction in sequence lengths as BPE vocabulary size increases.
IV-E Decoding methods

### v3

@@ Abstract
While recent research increasingly showcases the remarkable capabilities of Large Language Models (LLMs), it is equally crucial to examine their associated risks. Among these, privacy and security vulnerabilities are particularly concerning, posing significant ethical and legal challenges. At the heart of these vulnerabilities lies memorization, which refers to a model’s tendency to store and reproduce phrases from its training data. This phenomenon has been shown to be a fundamental source of various privacy and security issues related to LLMs.
In this paper, we provide a taxonomy of the literature on LLM memorization, structured around three dimensions: granularity, retrievability, and desirability. Building on this framework, we first examine the metrics and methods used to quantify memorization, highlighting how different approaches capture distinct aspects of the phenomenon. This foundation then supports an analysis of the factors that drive memorization and the strategies developed to mitigate its undesirable effects. Finally, we identify promising directions for future research, including methods to balance privacy and performance, as well as the study of memorization in specialized LLM contexts such as conversational agents, retrieval-augmented generation, and diffusion language models.
Given the rapid pace of research in this field, we also maintain a dedicated repository of the references discussed in this survey [^1] https://github.com/alistvt/undesirable-llm-memorization, which will be regularly updated to reflect the latest developments.
@@ I Introduction
Despite these impressive advancements of LLMs, researchers have shown that there are certain problems with these models, including hallucination [58], bias [141], and privacy and security vulnerabilities [156]. In the context of data privacy, memorization is one of the core sources of concern. Memorization in LLM refers to the model’s tendency to store and reproduce exact phrases or passages from the training data rather than generating novel or generalized outputs. While memorization can be advantageous in knowledge-intensive benchmarks, such as factual recall tasks or domain-specific question answering [112], it also introduces ethical and legal challenges: models may inadvertently reveal sensitive or private information included in their training data, posing significant privacy and security risks [20, 135]. In addition, the ability of LLMs to repeat verbatim copyrighted or proprietary text from their training data is closely related to memorization due to which many issues related to intellectual property infringement arise[6, 51].
These challenges motivate the need to further explore memorization as a phenomenon in LLMs to effectively tackle the associated challenges. In this paper, we provide an overview of aspects related to memorization, emphasizing the need for further exploration of this topic.
@@ I-B Data selection
The selection of works included in this survey was guided by the goal of capturing a comprehensive view of the literature surrounding memorization in LLMs, with a particular focus on its undesirable aspects. However, to provide a well-rounded taxonomy and to contextualize these challenges, we also include works that examine memorization in its desirable forms, such as generalization and knowledge retention, as they offer complementary insights into the mechanisms at play.
@@
The process of identifying relevant papers was iterative, beginning with the most cited early studies by recognized researchers in the field (14 papers). These foundational and influential works were selected as they represent key milestones that have shaped current understandings of memorization in LLMs. Then we included the papers drawn from keyword searches in article repositories. We used Google Scholar [^2] https://https://scholar.google.com/ and arXiv [^3] https://arxiv.org/, as they are predominantly used by researchers in the computer science and computational linguistics domains [138]. For Google Scholar we restricted our keyword search to the papers that include the term “Memorization” in their title and include “Language Model” in their body (memorization is also used as a term in biology). For the scanning of arXiv, we collected the papers that included both of the words “Memorization” and “Privacy” in their abstract and “language model” in their body. For inclusion, we selected articles published before December 2025.
@@
Figure 1: Our data selection process: Starting with 14 foundational papers as the seed, our literature search expanded through keyword search in Google Scholar (460 papers) and arXiv (185 papers), followed by manual filtering, reference chaining, and exploration of adjacent areas, concluding in a final selection of 120 papers.
@@ I-C Paper organization
Figure 2: Our categorization of the existing work on LLM memorization
@@ II Spectrum of Memorization
In this section, we provide an overview of the available definitions for memorization in three dimensions and categorize the existing work according to these axes. The overall summary of our findings is presented in Table III along with the other aspects.
@@ II-A Granularity of memorization
This definition is more strict and does not capture the memorization under different forms of sampling methods. It rather considers only the greedy sampling [see Section IV-C2].
@@
Throughout the rest of this paper, we use the word memorization interchangeably with the verbatim memorization (Definition Verbatim Memorization), as it is the one that is mostly associated with the undesirable [^4] Desirability is discussed in section II-C aspects of memorization.
@@ II-B Retrievability
[Extractable Memorization] Let $LM(p)$ denote the output of a language model when prompted with $p$. An example $s$ from the training set $S$ is extractably memorized if an adversary (without access to $S$) can construct a prompt $p$ that makes the model produce $s$ (that is, $LM(p)=s$).
@@
[Discoverable memorization] Let $LM(p)$ denote the output of a language model when prompted with $p$. For a context $c=(p,s)$ from the training set $S$, we say that $s$ is discoverably memorized if $LM(p)=s$.
@@
[k-discoverable memorization] For an example $c=(p,s)$ from the training set, string $s$ is said to be k-discoverable if $s$ is discoverable and $p$ is consisted of $k$ tokens [8].
@@
Hayes et al. [49] introduce the notion of (n,p)-discoverable extraction, which characterizes when an adversary can obtain a specific sample by querying the model n times with probability at least p, under defined sampling schemes. This offers a more operational and practice-oriented framework for assessing extraction risk.
@@ II-C Desirability
  • Copyright violation: Memorization of copyrighted material may lead to legal challenges, especially if the model can reproduce substantial portions of protected works [51, 97, 39, 27].
@@
  • Performance degradation on specific contexts: When models lean on memorized content, they may produce incorrect outputs on tasks that require reasoning, or generate invalid information due to reliance on outdated memorized knowledge [32, 114, 54].
@@
  • Transparency and auditing: When models retain recognizable traces of their training data, users and auditors can directly test whether specific information was included in the training corpus. This observability enables practical checks on data provenance, supports accountability, and offers a way for individuals to verify whether their own data may have been used, leading to a careful selection of training data by the model owners.
  • Knowledge retention: LLMs are often deliberately designed to retain facts, concepts, and general knowledge to enhance their performance in tasks like question answering and information retrieval [26, 86].
@@
  • Alignment goals: Intentional memorization can be utilized in the AI alignment phase to inject desired behaviors and values into the model [103].
Remark.
While the distinction between undesirable and desirable memorization is conceptually useful, the practical implications are more nuanced. Many effects of memorization depend on who is evaluating the model (e.g., model developers, end-users, data subjects, copyright holders) and in which context the memorized content appears (e.g., bias mitigation versus alignment, knowledge retention versus privacy). A behavior that benefits general users can simultaneously disadvantage copyright holders: memorization that enables auditing and data-provenance checks for users may expose sensitive information from data subjects; memorization that stabilizes alignment behaviors may also reinforce existing biases; memorization that improves factual recall may increase the risk of generation of false information due to out-datedness. In this sense, even when memorization provides functional advantages, it always carries an undesirable risk.
@@ III-A String match
Measuring verbatim memorization (Section Definition) with the methods discussed in the retrievability section (Section II-B), requires exhaustively interacting with the model by inputting different prompts and comparing the model output to the training data. Then the attack success rate is measured by dividing the portion of memorized text by the size of the training data. Since there are infinite combinations of tokens that one can feed the model, this method usually falls short in providing the accurate amount of memorization, rather it can provide a good approximation on memorization lower bound [19, 98, 49]. As the exact match metric is too sensitive to small perturbations in the generation, approximate match (Section Definition) has been introduced [61] to capture more memorized samples by bypassing negligible changes in the outputs.
@@ III-B Exposure metric
Definition.
[Exposure] Given a canary $s[r]$, a model with parameters $\theta$, and the randomness space R, the exposure of $s[r]$ is [19]:
@@ III-C Inference attacks
These attacks aim to determine whether a specific data point was part of the model’s training set. Shokri et al. [131] introduced this concept for machine learning models, and it has since been adapted for language models. Doing MIA on LLMs typically involves determining the model’s confidence in a given text and using it as an indicator of whether the text was part of the training data. This is often done using the log-likelihood of the sequence tokens [91, 126, 128, 63, 146, 136, 145, 150, 153, 123, 163].
@@
These attacks attempt to extract specific pieces of information from the model that were present in its training data. Different works have demonstrated the feasibility of such attacks on language models, showing that private information could be extracted through prompting with different strategies [62, 98, 108, 19, 87, 145, 49].
@@ III-D Counterfactuality
Previous work analyzed the memorization of large language models on sensitive information (e.g. phone numbers) in the training data [20] or synthetically injected ‘canaries’ [18, 52]. However, not all the memorized texts are equally interesting. Zhang et al. [162] propose another measure of memorization which is counterfactual memorization. The idea is to see how the presence or absence of a sample of the dataset, affects the performance of the model on the same sample. This measure has similarities with the definition of differential privacy. [^7] Differential privacy as a mitigation strategy will be discussed in Section V-B1.
@@
III-E Heuristic methods
Schwarzschild et al. [125] introduce the Adversarial Compression Ratio (ACR) as a novel metric to assess memorization in large language models (LLMs). ACR evaluates whether a string from the training data is memorized by determining if it can be elicited with a significantly shorter adversarial prompt. The method uses the GCG algorithm [167] to find the most compressed prompt. This approach aligns conceptually with Kolmogorov complexity [^8] Kolmogorov complexity measures the shortest possible set of instructions (or program) needed for a computer (Turing machine) to generate a specific output., as it measures the minimum description length of a string but adapts the concept for practical use in LLMs by focusing on adversarial prompts. Similarly, Feng et al. [38] utilize soft token sparsity as a measure of memorization, operating on the premise that LLMs generate memorized content with higher confidence and determinism.
@@ IV Influencing factors and dynamics of memorization
Understanding the factors that influence memorization in LLMs is a crucial to overcome the undesirabilities of this phenomenon. This section explores various aspects that affect memorization, organized into model-related aspects, training pipeline influences, and inference-time causes.
IV-A Model-related factors & dynamics
IV-A1 Model capacity
@@
IV-A2 Tokenization
@@
Kharitonov et al. [70] explore the impact of the tokenizer on memorization. They experiment with the size of the sub-word vocabulary learned through Byte-Pair Encoding (BPE) and demonstrate that increasing the sub-word vocabulary significantly affects the model’s ability and inclination to memorize training data. Furthermore, models with larger vocabularies are more likely to reproduce training data when given specific prompts. The authors suggest that this stems from the reduction in sequence lengths as BPE vocabulary size increases.
IV-A3 Explainability and interpretability
A growing body of interpretability work aims to uncover how LLMs memorization is represented, triggered, and propagated. Huang et al. [56] study verbatim memorization in LLMs using controlled pre-training with injected sequences. They find that memorization requires significant repetition, increases in later checkpoints, and is tied to distributed model states and general language modeling capabilities. Their stress tests show unlearning methods often fail to remove memorized information without degrading model performance, highlighting the difficulty of isolating memorization. Haviv et al. [48] propose a framework to probe how memorized sequences are recalled in transformers, showing that memory recall follows a two-step process: early layers promote the correct token in the output distribution, while upper layers amplify confidence. They find that memorized information is primarily stored and retrieved in early layers. Building on this layer-wise perspective, Dankers and Titov [28] investigate where memorization occurs across model layers, demonstrating that memorization is a gradual and task-dependent process rather than localized to specific layers. Using centroid analysis, they show that deeper layers contribute more to memorization when models generalize well to new data. Complementing these findings, Stoehr et al. [137] show that memorization in LLMs, while distributed across layers, is driven by distinct gradients in lower layers and influenced by a low-layer attention head focusing on rare tokens. Perturbation analysis reveals that distinctive tokens in a prefix can corrupt entire continuations, and memorized sequences are harder to unlearn and more robust to corruption than non-memorized ones.
Beyond layer-level analysis, other work explores how memorization manifests in the model’s internal representations. Chen et al. [24] identify clustering of sentences with different memorization scores in the embedding space and observe an inverse boundary effect in entropy distributions for memorized and unmemorized sequences. They also demonstrate that hidden states of LLMs can be used to predict unmemorized tokens, shedding light on the dynamics of memorization within the model.
More mechanistic perspectives further refine this picture. Huang et al. [57] show that memorization and generalization in LLMs are encoded in largely distinct subsets of neurons. By intervening on these neurons at inference, they demonstrate that a model’s tendency to memorize or generalize can be selectively amplified or suppressed without substantially affecting overall performance. Lasy et al. [76] show that verbatim memorization in LLMs is driven by specific neural circuits, with distinct subgraphs responsible for initiating versus sustaining memorized sequences, revealing mechanistic pathways underlying exact data recall.
Together, these findings offer ongoing insights into the mechanisms of memorization, enhancing interpretability and guiding future research.
IV-B Training pipeline
IV-B1 Training data characteristics
The nature of the training data heavily influences memorization. Lee et al. [77] develop tools to deduplicate training data and show that models trained on deduplicated data would produce memorized text ten times less frequently. Kandpal et al. [66] show that a sequence appearing 10 times in the training data is, on average, generated approximately 1000 times more frequently than a sequence that appears only once. These findings emphasize the central role of duplication in driving memorization.
A study by Tirumala et al. [140] on the memorization of different parts of speech in a text reveals that nouns and numbers are memorized significantly faster than other parts of speech, likely because they serve as unique identifiers for specific samples. Prashanth et al. [114] categorize memorization into recitation (memorization of highly duplicated sequences), reconstruction (generation using learned templates or patterns), and recollection (memorization of rare, non-template sequences). They argue that memorization is a multi-faceted phenomenon influenced by duplication, template patterns, and rarity, requiring nuanced analysis.
As simple sequences, such as repeated patterns or numbers, are easily memorized by models but often lack substantive content or sensitive information, distinguishing the memorization of these trivial sequences from more complex and meaningful ones would be essential. Duan et al. [33] observed that data with lower z-complexity [^9] Z-complexity is a way to measure how much temporary memory (workspace) a computer needs to generate a specific output using the shortest possible instructions. Unlike Kolmogorov complexity, which only looks at the length of the instructions, Z-complexity also considers the memory used during the process. leads to faster decreases in training loss, as more compressible patterns are memorized more quickly. They further demonstrated that strings of varying complexity exhibit distinct memorization curves, with lower-complexity strings being memorized more easily even for smaller repeats, following a log-linear relationship in memorization probability. This is indirectly supported by the study of Arnold [4] on the memorization as a function of “Intrinsic Dimension” [^10] Intrinsic dimension is the minimum number of variables (or degrees of freedom) needed to describe the essential structure of a dataset or a model’s representations., observing a negative correlation for the low-duplicated data points group.
Across these perspectives, the evidence converges: duplication, content distinctiveness, and structural simplicity jointly determine how easily different data types are memorized.
IV-B2 Training process dynamics
Beyond data properties, the temporal dynamics of training also plays a critical role in shaping when and how memorization emerges. Zhang et al. [162], Kandpal et al. [66] show that memorization grows consistently with the number of training epochs, which makes sense, as more epochs push the model to potential overfitting. Jagielski et al. [64] demonstrate that examples seen during the earlier stages of training are less prone to memorization and rather they are forgotten over time. These findings indicate that memorization increases with more training, while early-seen examples being more likely to be forgotten. Leybzon and Kervadec [79] show higher memorization rates happens early and late in training, with lower rates mid-training. Therefore, suggesting that placing sensitive data in the middle stages of training could reduce its vulnerability to extraction attacks.
IV-B3 Forgetting mechanisms
Kirkpatrick et al. [73] initially introduced “catastrophic forgetting” in the context of neural networks and continual learning. They propose a method to protect important model weights to retain knowledge. This approach has been effective in maintaining performance on older tasks, even after long periods of non-use. While this early work focused on task retention, later studies explore forgetting within pretraining settings.
Tirumala et al. [140] observe the forgetting mechanisms of a special batch through the learning process and show that it follows an exponential degradation, reaching a constant value baseline. They show that the mentioned baseline scales with the model size. These results suggest that forgetting is structured rather than arbitrary, and that larger models may retain certain information more persistently (section IV-A1).
Jagielski et al. [64] address the dual phenomena of memorization and forgetting in LLMs through stronger privacy attacks and several strategies for measuring the worst-case forgetting of the training examples. The study introduces a method to measure to what extent models forget specific training data, highlighting that standard image, speech, and language models do indeed forget examples over time, though non-convex models might retain data indefinitely in the worst case. The findings suggest that examples from early training phases, such as those used in pre-training large models, might enjoy privacy benefits but could disadvantage examples encountered later in training.
Overall, these studies illustrate that forgetting is shaped by both the training timeline and the model’s internal mechanisms, and can sometimes be leveraged to balance memorization, generalization, and privacy.
IV-B4 Fine-tuning and transfer learning
IV-C Inference-time factors
IV-C1 Input and prompting strategies
Prompting methods can strongly influence the generation of memorized content at the inference stage. Carlini et al. [20], McCoy et al. [92], Kandpal et al. [66] show that longer prompts increase the likelihood of triggering memorized sequences, making it easier for language models to regurgitate training data. Moreover, methods like prefix tuning [81] and prompt engineering have been employed to maximize memorization. Ozdayi et al. [105] introduce a novel approach using prompt tuning to control memorized content extraction rates in LLMs. Wang et al. [147] introduce a dynamic, prefix-dependent soft prompt approach to elicit more memorization. By generating soft prompts based on input variations, this method outperforms previous techniques in extracting memorized content across both text and code generation tasks. Kassem et al. [67] investigate how instruction-tuning influences memorization in LLMs, showing that instruction-tuned models can reveal as much or more training data as base models. They propose a black-box prompt optimization method where an attacker LLM generates instruction-based prompts, achieving higher memorization levels than direct prompting approaches. Weller et al. [151] propose ‘according-to’ prompting, a technique that directs LLMs to ground responses in previously observed text.
IV-C2 Decoding methods